OpenStudy (anonymous):

ok, just need verification on this one, jack can mow his lawn in 3 hours, sara can mow it in 6, how many hours would it take working together? i got 2. is this right?

OpenStudy (saifoo.khan):

wrong! :(

OpenStudy (anonymous):

whaaaaaa!! help!

OpenStudy (saifoo.khan):

Call 911 asap for help!

OpenStudy (anonymous):

lol

OpenStudy (saifoo.khan):

\[\frac{1}{3} + \frac 16 = \frac 1x\]

OpenStudy (saifoo.khan):

solve for x.

OpenStudy (anonymous):

jack working for 1 hour can mow 1/3 of his lawn at same time sara mows 1/6 that should help ogether they have mown 1/2 the lawn

OpenStudy (anonymous):

shouldn't it be x/3 + x/6 = 1? simplify to 3x+6x=18

OpenStudy (anonymous):

9x=18 x=2?

OpenStudy (saifoo.khan):

Nope! :(

OpenStudy (jamesj):

What Jimmy said is a very good way to think about it: think about rates of doing things. Saifoo is saying the same thing as well, but perhaps slightly less obvious.

OpenStudy (anonymous):

so 2 hours is right then, if according to jimmy it takes 1 hour for the to mow 1/2 the lawn together, it will take them 2 hours to mow the whole thing!